In a period of 1.5 s, 5.0 1023 nitrogen molecules strike a wall of area 7.2 cm2. If the molecules move at 310 m/s and strike the wall head on in a perfectly elastic collision, find the pressure exerted on the wall. (The mass of one N2 molecule is 4.68 10-26 kg.)
Answer #1

collision is perfectly elastic hence after collison molecules will move in opposite direction with same speed.

hence change in momentum for 1 molecule = 4.68 x 10^-26 x (310 + 310) = 2.902 x 10^-23

total momentum transferred in 1.5 sec, p = 5 x 10^23 x 2.902 x 10^-23 = 14.51 kg m/s

Force = p/t = 14.51 / 1.5 = 9.67 N

and pressure, P = F/A = 9.67 / (7.2 x 10^-4) = 13433.33 Pa